# A Prompt Learning Framework for Source Code Summarization

### v1

@@ Abstract.
(Source) code summarization is the task of automatically generating natural language summaries (also called comments) for given code snippets. Such summaries play a key role in helping developers understand and maintain source code. Recently, with the successful application of large language models (LLMs) in numerous fields, software engineering researchers have also attempted to adapt LLMs to solve code summarization tasks. The main adaptation schemes include instruction prompting and task-oriented fine-tuning. However, instruction prompting involves designing crafted prompts for zero-shot learning or selecting appropriate samples for few-shot learning and requires users to have professional domain knowledge, while task-oriented fine-tuning requires high training costs.
@@
In this paper, we propose a novel prompt learning framework for code summarization called PromptCS. It no longer requires users to rack their brains to design effective prompts. Instead, PromptCS trains a prompt agent that can generate continuous prompts to unleash the potential for LLMs in code summarization. Compared to the human-written discrete prompt, the continuous prompts are produced under the guidance of LLMs and are therefore easier to understand by LLMs. PromptCS is non-invasive to LLMs and freezes the parameters of LLMs when training the prompt agent, which can greatly reduce the requirements for training resources. We evaluate the effectiveness of PromptCS on the CodeSearchNet dataset involving multiple programming languages. Experimental results show that PromptCS significantly outperforms instruction prompting schemes (including zero-shot learning and few-shot learning) on all four widely used metrics, including BLEU, METEOR, ROUGH-L, and SentenceBERT, and is comparable to the task-oriented fine-tuning scheme. In some base LLMs, e.g., CodeGen-Multi-2B and StarCoderBase-1B and -3B, PromptCS even outperforms the task-oriented fine-tuning scheme. More importantly, the training efficiency of PromptCS is faster than the task-oriented fine-tuning scheme, with a more pronounced advantage on larger LLMs. The results of the human evaluation demonstrate that PromptCS can generate more good summaries compared to baselines.
@@
source code summarization, large language model, prompt learning
copyright: acmcopyrightjournalyear: 2023journal: TOSEMjournalvolume: 0journalnumber: 0article: 1publicationmonth: 0ccs: Software and its engineering Software maintenance tools
1. Introduction
@@
Source code comments play a critical role in facilitating program comprehension (Tenny, 1988; Stapleton et al., 2020; Sun et al., 2023a) and software maintenance (de Souza et al., 2005; Sun et al., 2023c). However, existing research (Hu et al., 2018b, 2022b) demonstrates that lack of high-quality code comments is a common problem in the software industry. In addition, comments are often absent, unmatched, and outdated during software evolution (de Souza et al., 2005). These practical problems drive the research on source code summarization. Source code summarization (code summarization for short) is the task of automatically generating code summaries (i.e., comments). Over the past decade, it has always been one of the research hotspots in the field of software engineering (Haiduc et al., 2010a; Eddy et al., 2013; Hu et al., 2018a; Zhang et al., 2020; Stapleton et al., 2020; Shi et al., 2022; Bansal et al., 2023; Wang et al., 2023a; Sun et al., 2023c; Tian et al., 2023).
@@
Recently, with the success of large language models (LLMs) in natural language processing (NLP) (Du et al., 2022; Qin et al., 2023), an increasing number of software engineering (SE) researchers have started integrating them into the resolution process of various SE tasks (Chen et al., 2021; Fan et al., 2023; Hou et al., 2023). Similar to LLMs for NLP (e.g., ChatGPT (OpenAI, 2022) and LLaMA (Touvron et al., 2023)) tasks, there are many LLMs of code for SE tasks, e.g., Codex (OpenAI, 2023a), StarCoder (Li et al., 2023), CodeGen (Nijkamp et al., 2023), and PolyCoder (Xu et al., 2022). In this paper, we pay more attention to the application of LLMs on code summarization tasks. Existing studies demonstrate that LLMs trained on massive corpora of texts have shown their particularly exciting ability to perform new tasks from textual instructions (i.e., in zero-shot learning setting) or from a few examples (i.e., in few-shot learning setting) (Brown et al., 2020; Ahmed and Devanbu, 2022b). Therefore, there have been several recent studies investigating the effectiveness of instruction prompting with zero-shot and few-shot learning in adapting LLMs to code summarization tasks. For example, Sun et al. (Sun et al., 2023c) evaluate ChatGPT’s performance on zero-shot code summarization tasks. They design several heuristic questions/instructions to collect the feedback of ChatGPT, thereby finding an appropriate prompt to guide ChatGPT to generate in-distribution code summaries. However, their experimental results on the large-scale CSN-Python dataset show that this prompt fails to induce ChatGPT to generate satisfactory summaries. We also try this prompt on the CSN-Java dataset and get similar results, detailed in Section 3. These results suggest that crafting effective prompts manually is a challenging task, demanding users to possess not only specialized domain knowledge but also a profound understanding of the LLMs being utilized (Wang et al., 2023b; Sun et al., 2023c). Ahmed et al. (Ahmed and Devanbu, 2022b) investigate the effectiveness of the instruction prompting with few-shot learning in adapting LLMs to code summarization tasks. To discover the appropriate example number for few-shot learning, they try multiple sets of samples with different numbers (including 5, 10, and 15) on a small-scale test dataset. Then, they utilize 10 samples to conduct instruction prompting on a commercial LLM Codex and find it can outperform fine-tuned foundation models (e.g., CodeT5 (Wang et al., 2021)). We utilize the same 10 samples to conduct instruction prompting on open-source LLMs (e.g., StarCoderBase-3B) and find that their performance is still much worse than fine-tuned LLMs, detailed in Section 3 and Section 5.2.1. This phenomenon indicates that for few-shot learning, which samples to choose and how many samples to choose require users to make decisions based on professional knowledge and continuous trials.
@@
Another straightforward scheme to adapt LLMs to better accomplish downstream tasks is through task-oriented fine-tuning (Jin et al., 2023; Xu et al., 2021). Task-oriented fine-tuning updates the weights of LLMs by training on thousands of supervised labels specific to the desired task. The main advantage of fine-tuning is strong performance on the task benchmark (Brown et al., 2020; Radford et al., 2019; Wei et al., 2022). For example, Jin et al. (Jin et al., 2023) fine-tune Codex with 12 billion (B) on supervised bug-fix data. Although fine-tuning significantly enhances the model’s ability in bug fixing, it comes with a high training cost. For instance, to conduct full model fine-tuning (updating all weights of Codex), Jin et al. build a training environment consisting of 64 32GB V100 GPUs, which is beyond the reach of many research teams. In this paper, we also experiment with the task-oriented fine-tuning scheme on code summarization tasks, which similarly incurred a high training cost, detailed in Section 5.2.1. As the size of LLMs continues to grow, the cost of fine-tuning can be quite substantial.
@@
In this paper, we propose a novel prompt learning framework for code summarization, called PromptCS. The key feature of PromptCS is its ability to free developers from the need to manually design intricate prompts. Specifically, PromptCS devises two collaborating components to achieve this feature, including a prompt agent and an LLM. The prompt agent is responsible for generating the continuous prompt that can induce the capacity of the LLM to perform code summarization tasks. The core of the prompt agent is a deep learning (DL) based prompt encoder, which takes a pseudo prompt consisting of $n$ learnable tokens as input and produces a prompt embedding (i.e., continuous prompt). The prompt agent is trained under the guidance of the LLM. Therefore, the well-trained prompt agent can produce a continuous prompt that is more suitable for the LLM than the human-written discrete prompt. More importantly, unlike the task-oriented fine-tuning scheme that changes the parameters of the LLMs, PromptCS is non-invasive to LLMs. PromptCS freezes the parameters of LLMs during the training process and only updates the parameters of the prompt agent, which can greatly reduce the requirements for training resources.
@@
  • We propose a novel prompt learning framework for code summarization called PromptCS. PromptCS is able to generate high-quality summaries for input code snippets. In addition, PromptCS is a general framework and can be combined with multiple LLMs (see experimental results in Section 5.2.3). We open source the code of PromptCS (Sun et al., 2023b) to facilitate future research and application.
@@
  • We conduct a qualitative human evaluation to evaluate the summaries generated by PromptCS and baselines (including instruct prompting schemes with zero-shot and few-shot learning and the task-oriented fine-tuning scheme. The statistical results of human scores show that compared with baselines, the summaries generated by PromptCS are more similar to the ground-truth summaries (detailed in Section 5.2.6).
2. background
2.1. Code Summarization
Figure 1. An example of a code snippet and its summary
@@ 2.1. Code Summarization
Code summarization is the task of automatically generating natural language summaries for code snippets given by the developer. Such summaries are also called comments and explain the functionality of the code snippet (Moore et al., 2019; Hu et al., 2020, 2022b). Figure 1 shows an example. The code snippet $c_{1}$ in Figure 1(a) is provided by the developer. The summary “destroy kills the connection and drops it from the connection pool. destroy should only occur if the connection state machine has failed.” in Figure 1(b) is a summary that satisfies the developer’s requirement. The research on code summarization can be traced back to as early as 2010 when Sonia Haiduc et al. (Haiduc et al., 2010a) introduced automated text summarization technology to summarize source code to support program comprehension. Later on, following the significant success of neural machine translation (NMT) research in the field of NLP (Cho et al., 2014a; Bahdanau et al., 2015), a large number of researchers migrate its underlying DL-based encoder-decoder architecture to code summarization tasks (Zhang et al., 2020; Gros et al., 2020; Ahmad et al., 2020; Shi et al., 2022; Sun et al., 2023a; Wang et al., 2023a; Bansal et al., 2023). In the past ten years, code summarization has always been one of the hot research directions in the field of software engineering. In this paper, similar to (Ahmed and Devanbu, 2022b; Sun et al., 2023c; Tian et al., 2023), we focus on adapting LLMs for code summarization tasks. More details of related works are discussed in Section 7.1.
2.2. Large Language Model
@@
Existing research (Kaplan et al., 2020) demonstrates that scaling pre-trained language models (PLMs), e.g., increasing model size, can enhance model capacity for solving downstream tasks. To distinguish the early small-scale PLMs, more people are now willing to refer to the current generative PLMs with very large-scale parameters (often in the billions) as large language models (LLMs). When utilizing LLMs to accomplish downstream tasks, three components play a crucial role: a tokenizer, an input embedding layer, and module blocks. The tokenizer is responsible for converting the input text into the index representation that the model can understand, serving as a bridge between the input text and the model. The primary function of the input embedding layers is to transform the index representation into vector representations (also called embeddings). This process may involve techniques like word embeddings (Pennington et al., 2014) and positional encodings (Vaswani et al., 2017) to capture vocabulary and positional information. These techniques enhance the input embedding layer’s ability to transform discrete data into meaningful continuous embeddings, facilitating subsequent processing. In the module blocks, the neural network structure is carefully designed to efficiently capture the complex relationships within embeddings. These blocks typically consist of multiple layers, with each layer responsible for processing different levels of features and abstract representations.
@@
Currently, LLMs can be mainly divided into three categories: auto-regressive language models, masked language models, and encoder-decoder language models (Xu et al., 2022). The auto-regressive language models generate sequences based on the outputs from previous time steps. They are powerful for modeling the probability of sequences. In the code summarization task, the model recursively predicts each word to generate a code summary. The currently popular LLMs all belong to this category, such as GPT series (OpenAI, 2023b), Codex (OpenAI, 2023a), StarCoder (Li et al., 2023), CodeGen (Nijkamp et al., 2023), PolyCoder (Xu et al., 2022), and CodeLlama (Rozière et al., 2023). The masked language model can transform a given sequence into effective representations. They are trained using masked language modeling, a popular bidirectional objective function that aims to predict masked text pieces based on the surrounding context. Most of the early PLMs fall into this category, such as CodeBERT (Feng et al., 2020), UniXCoder (Guo et al., 2022), and GraphCodeBERT (Guo et al., 2021). The encoder-decoder language models consist of a structure with two main components, i.e., encoder and decoder. The encoder is responsible for transforming the input sequence into an intermediate representation, while the decoder utilizes this intermediate representation to generate the output sequence. CodeT5 (Wang et al., 2021) and PLBART (Ahmad et al., 2021) are examples of such models in code. In recent times, the impact of LLMs has been truly transformative, revolutionizing both academic research and various industrial applications. In this paper, we focus on the application of autoregressive LLMs in software engineering, particularly in code summarization.
3. Motivation
In this section, we elaborate on the shortcomings of the latest schemes for adapting LLMs to code summary tasks as the motivation for us to propose a new scheme. To facilitate understanding the effect of each scheme, we perform the task-oriented fine-tuning scheme on an LLM called StarCoderBase-3B on the CSN-Java dataset and use its results as a reference. Existing works (Radford et al., 2019; Wei et al., 2022; Weyssow et al., 2023) demonstrate that LLMs often achieve peak effectiveness and show their full potential when subjected to fine-tuning. Detailed results are shown in row 3 of Table 1.
Table 1. Existing LLM-based methods on code summarization.
Method Base Model Model Size CSN-Java
BLEU METEOR ROUGE-L
Task-oriented Fin-tuning StarCoderBase 3B 20.43 14.68 39.51
Instruction Prompting (zero-shot) ChatGPT UNK1 9.57 14.89 20.75
Instruction Prompting (few-shot) StarCoderBase 3B 15.12 11.64 29.84
PromptCS StarCoderBase 3B 20.87 14.50 40.23
  • 1
    OpenAI has not disclosed the specific parameter scale of ChatGPT.
As mentioned in Section 1, currently, there are several schemes available for researchers to adapt LLMs to code summarization tasks. Among them, the most prominent approach is instruction prompting with zero-shot/few-shot learning (Ahmed and Devanbu, 2022b; Sun et al., 2023c; Bhattacharya et al., 2023). For example, Sun et al. (Sun et al., 2023c) adopt instruction prompting with zero-shot learning to induce ChatGPT to complete code summarization tasks. To gain a suitable prompt, they try out various prompts by adjusting details such as language choice and grammatical structure. After many complicated attempts, they finally find an appropriate prompt that can be used to guide ChatGPT to generate in-distribution comments, that is “Please generate a short comment in one sentence for the following function:¡code¿”. We utilize this prompt to guide ChatGPT in generating summaries for code snippets in the CSN-Java dataset, and results are shown in row 4 of Table 1. From Table 1, it can be observed that the results of ChatGPT are similar to those on the CSN-Python dataset shown in their paper (Sun et al., 2023c) (10.28 in BLEU, 14.40 in METEOR, and 20.81 in ROUGE-L), which are both unsatisfactory compared to the task-oriented fine-tuning results. It indicates that even professional researchers in the field of code summarization have difficulty designing good prompts.
Apart from the work (Sun et al., 2023c), Ahmed et al. (Ahmed and Devanbu, 2022b) investigate the effectiveness of instruction prompting with few-shot learning for code summarization tasks. They conduct few-shot learning on a commercial LLM of code called Codex (OpenAI, 2023a), and find it can significantly outperform a fine-tuned model trained with thousands of samples with just ten samples. We experiment with the same few-shot learning samples on multiple open-source LLMs of code. Row 4 of Table 1 shows the experimental results on StarCoderBase-3B. Observe that although compared with the instruction prompting with zero-shot learning, the instruction prompting with few-shot learning can improve the performance of the LLM on code summarization in terms of BLEU and ROUGE-L, it is still much worse than the task-oriented fine-tuning results. It means that for the instruction prompting with few-shot learning scheme, selecting appropriate code summarization samples for few-shot learning is also a complex task. Deciding which and how many samples to select requires expertise in the field and trial and error.
Task-oriented fine-tuning is also a straightforward idea to adapt LLMs to downstream tasks. However, unlike instruction prompting schemes that are non-invasive to LLMs, the fine-tuning process involves updating and storing the parameters of LLMs. In practical applications, researchers and practitioners may need to invest significant resources to complete the fine-tuning process to ensure the fine-tuned LLMs perform well in a specific domain or downstream task. Task-oriented fine-tuning is very time-consuming, especially when there are few training resources available. For instance, our experiments show that it takes more than 200 hours to fine-tune the StarCoder-7B model only for one epoch on a 1-card A800GPU device, detailed in Section 5.2.1. As LLMs continue to grow in scale, the training time and computational costs associated with task-oriented fine-tuning become increasingly challenging to tolerate. In addition, it should be noted that fine-tuning changes in the parameters of LLMs may affect their performance on other downstream tasks.
Our solution. We propose a novel prompt learning framework PromptCS for adapting LLMs to code summarization tasks. Unlike instruction prompting techniques that require complex prompt design or carefully selected samples for few-shot learning, PromptCS trains a prompt agent that can generate continuous prompts inducing LLMs to generate high-quality summaries for given code snippets. Unlike the fine-tuning technique that changes the parameters of the LLMs, PromptCS is also non-invasive to LLMs, freezes the parameters of LLMs, and only updates the parameters of the prompt agent, greatly reducing the requirements for training resources. The last row of Table 1 shows the results of PromptCS on the CSN-Java dataset. Observe that PromptCS significantly outperforms the instruction prompting schemes (including zero-shot learning and few-shot learning) on all three metrics, even surpassing task-oriented fine-tuning on BLEU and ROUGE-L. In Section 5.2.1, we empirically demonstrate that PromptCS outperforms three state-of-the-art baselines in improving metric scores.
4. Methodology
4.1. Overview

### v2

@@ Abstract
(Source) code summarization is the task of automatically generating natural language summaries (also called comments) for given code snippets. Recently, with the successful application of large language models (LLMs) in numerous fields, software engineering researchers have also attempted to adapt LLMs to solve code summarization tasks. The main adaptation schemes include instruction prompting, task-oriented (full-parameter) fine-tuning, and parameter-efficient fine-tuning (PEFT). However, instruction prompting involves designing crafted prompts and requires users to have professional domain knowledge, while task-oriented fine-tuning requires high training costs, and effective, tailored PEFT methods for code summarization are still lacking.
@@
In this paper, we propose an effective prompt learning framework for code summarization called PromptCS. It no longer requires users to rack their brains to design effective prompts. Instead, PromptCS trains a prompt agent that can generate continuous prompts to unleash the potential for LLMs in code summarization. Compared to the human-written discrete prompt, the continuous prompts are produced under the guidance of LLMs and are therefore easier to understand by LLMs. PromptCS is non-invasive to LLMs and freezes the parameters of LLMs when training the prompt agent, which can greatly reduce the requirements for training resources. We evaluate the effectiveness of PromptCS on the CodeSearchNet dataset involving multiple programming languages. Experimental results show that PromptCS significantly outperforms instruction prompting schemes (including zero-shot learning and few-shot learning) on all four widely used metrics, including BLEU, METEOR, ROUGE-L, and SentenceBERT, and is comparable to the task-oriented fine-tuning scheme. In some base LLMs, e.g., CodeGen-Multi-2B and StarCoderBase-1B and -3B, PromptCS even outperforms the task-oriented fine-tuning scheme. More importantly, the training efficiency of PromptCS is faster than the task-oriented fine-tuning scheme, with a more pronounced advantage on larger LLMs. The results of the human evaluation demonstrate that PromptCS can generate more good summaries compared to baselines.
keywords:
@@
source code summarization, large language model, prompt learning, AI for SE, parameter-efficient fine-tuning
1 Introduction
@@
Source code comments play a critical role in facilitating program comprehension [1, 2, 3] and software maintenance [4, 5]. However, existing research [6, 7] demonstrates that lack of high-quality code comments is a common problem in the software industry. In addition, comments are often absent, unmatched, and outdated during software evolution [4]. These practical problems drive the research on source code summarization. Source code summarization (code summarization for short) is the task of automatically generating code summaries (i.e., comments). Over the past decade, it has always been one of the research hotspots in the field of software engineering [8, 9, 10, 11, 2, 12, 13, 14, 5, 15].
@@
Recently, with the success of large language models (LLMs) in natural language processing (NLP) [16, 17], an increasing number of software engineering (SE) researchers have started integrating them into the resolution process of various SE tasks [18, 19, 20]. Similar to LLMs for NLP (e.g., ChatGPT [21] and LLaMA [22]) tasks, there are many LLMs of code for SE tasks, e.g., Codex [23], StarCoder [24], CodeGen [25], and PolyCoder [26]. In this paper, we pay more attention to the application of LLMs on code summarization tasks. Existing studies demonstrate that LLMs trained on massive corpora of texts have shown their particularly exciting ability to perform new tasks from textual instructions (i.e., in zero-shot learning setting) or from a few examples (i.e., in few-shot learning setting) [27, 28]. Therefore, there have been several recent studies investigating the effectiveness of instruction prompting with zero-shot and few-shot learning in adapting LLMs to code summarization tasks. For example, Sun et al. [5] evaluate ChatGPT’s performance on zero-shot code summarization tasks. They design several heuristic questions/instructions to collect the feedback of ChatGPT, thereby finding an appropriate prompt to guide ChatGPT to generate in-distribution code summaries. However, their experimental results on the large-scale CSN-Python dataset show that this prompt fails to induce ChatGPT to generate satisfactory summaries. We utilize their prompt to guide ChatGPT in generating summaries and the result is unsatisfactory compared to the result of task-oriented fine-tuning. It indicates that even professional researchers in the field of code summarization have difficulty designing good prompts. Ahmed et al. [28] investigate the effectiveness of the instruction prompting with few-shot learning in adapting LLMs to code summarization tasks. To discover the appropriate example number for few-shot learning, they try multiple sets of samples with different numbers (including 5, 10, and 15) on a small-scale test dataset. Then, they utilize 10 samples to conduct instruction prompting on Codex and find it can outperform fine-tuned foundation models (e.g., CodeT5 [29]). We utilize the same 10 samples to conduct instruction prompting on open-source LLMs (e.g., StarCoderBase-3B) and find that their performance is still much worse than fine-tuned LLMs, detailed in Section 4.2.1. This phenomenon indicates that for few-shot learning, which samples to choose and how many samples to choose require users to make decisions based on professional knowledge and continuous trial, as demonstrated by the work [30].
@@
Another straightforward scheme to adapt LLMs to better accomplish downstream tasks is task-oriented (full-parameter) fine-tuning [31, 32]. Task-oriented fine-tuning updates the weights of LLMs by training on thousands of supervised labels specific to the desired task. The main advantage of fine-tuning is strong performance on the task benchmark [27, 33, 34]. For example, Jin et al. [31] fine-tune Codex with 12 billion (B) on supervised bug-fix data. Although fine-tuning significantly enhances the model’s ability in bug fixing, it comes with a high training cost. For instance, to conduct full model fine-tuning (updating all weights of Codex), Jin et al. build a training environment consisting of 64 32GB V100 GPUs, which is beyond the reach of many research teams. In this paper, we also experiment with the task-oriented fine-tuning scheme on code summarization tasks, which similarly incurs a high training cost, detailed in Section 4.2.1. As the size of LLMs continues to grow, the cost of fine-tuning can be quite substantial. In recent years, researchers in NLP have proposed numerous parameter-efficient fine-tuning (PEFT) methods to reduce the cost of adapting LLMs, such as P-tuning [35], Prefix-tuning [36], and LoRA [37]. PEFT methods can achieve comparable performance to task-oriented fine-tuning by updating fewer parameters [38, 39, 35]. Inspired by these works, many SE researchers have also begun investigating the performance of PEFT methods for adapting LLMs to code-related tasks [40, 41, 42, 43], detailed in Section 6.2. Most of these works are empirical studies; although their experiments cover various PEFT methods and a range of code-related tasks, they lack in-depth analysis of the performance of specific PEFT methods and the factors influencing their effectiveness on particular code-related tasks. Consequently, effective, customized PEFT methods for code summarization tasks are still lacking.
@@
In this paper, we propose a prompt learning framework specifically for code summarization, called PromptCS. Like P-tuning and Prefix-tuning, the key feature of PromptCS is its ability to free developers from the need to manually design intricate prompts. Specifically, PromptCS devises two collaborating components to achieve this feature, including a prompt agent and an LLM. The prompt agent is responsible for generating the continuous prompt that can induce the capacity of the LLM to perform code summarization tasks. The core of the prompt agent is a deep learning (DL) based prompt encoder, which takes a pseudo prompt consisting of $n$ learnable tokens as input and produces a prompt embedding (i.e., continuous prompt). The prompt agent is trained under the guidance of the LLM. Therefore, the well-trained prompt agent can produce a continuous prompt that is more suitable for the LLM than the human-written discrete prompt. More importantly, unlike the task-oriented fine-tuning scheme that changes the parameters of the LLMs, PromptCS is non-invasive to LLMs. PromptCS freezes the parameters of LLMs during the training process and only updates the parameters of the prompt agent, which can greatly reduce the requirements for training resources. Additionally, the non-intrusive approach offers another advantage: multiple downstream tasks can utilize the same base LLM, requiring only task-specific prompt agents to be trained. These prompt agents do not alter the parameters of the LLM, preserving its performance on other tasks. When new downstream tasks arise, a new prompt agent can simply be trained without impacting the LLM’s effectiveness on existing tasks.
@@
  • We propose a novel prompt learning framework for code summarization called PromptCS. PromptCS is able to generate high-quality summaries for input code snippets. In addition, PromptCS is a general framework and can be combined with multiple LLMs (see experimental results in Section 4.2.3). We open source the code of PromptCS [44] to facilitate future research and application.
@@
  • We conduct a qualitative human evaluation to evaluate the summaries generated by PromptCS and baselines (including instruct prompting schemes with zero-shot and few-shot learning and the task-oriented fine-tuning scheme. The statistical results show that compared with baselines, the summaries generated by PromptCS are more similar to the ground-truth summaries (detailed in Section 4.2.6).
The rest of this paper is organized as follows. Section 2 describes the background of code summarization and large language model. Section 3 introduces the design of PromptCS. Section 4 presents the design of the experiments in detail and gives the details of experiment results and analysis. Section 5 presents some code summarization cases. Section 6 discusses the related work of this paper. We conclude the paper in Section 7.
2 Background
2.1 Code Summarization
Figure 1: An example of a code snippet and its summary
@@ 2.1 Code Summarization
Code summarization is the task of automatically generating natural language summaries for code snippets given by the developer. Such summaries are also called comments and explain the functionality of the code snippet [45, 46, 7]. Figure 1 shows an example. The code snippet $c_{1}$ in Figure 1(a) is provided by the developer. The summary “destroy kills the connection and drops it from the connection pool. destroy should only occur if the connection state machine has failed.” in Figure 1(b) is a summary that satisfies the developer’s requirement. The research on code summarization can be traced back to as early as 2010 when Sonia Haiduc et al. [8] introduced automated text summarization technology to summarize source code to support program comprehension. Later on, following the significant success of neural machine translation (NMT) research in the field of NLP [47, 48], a large number of researchers migrate its underlying DL-based encoder-decoder architecture to code summarization tasks [11, 49, 50, 12, 3, 14, 13]. In the past ten years, code summarization has always been one of the hot research directions in the field of software engineering. In this paper, similar to [28, 5, 15], we focus on adapting LLMs for code summarization tasks. More details of related works are discussed in Section 6.1.
2.2 Large Language Model
@@
Existing research [51] demonstrates that scaling pre-trained language models (PLMs), e.g., increasing model size, can enhance model capacity for solving downstream tasks. To distinguish the early small-scale PLMs, more people are now willing to refer to the current generative PLMs with very large-scale parameters (often in the billions) as large language models (LLMs). When utilizing LLMs to accomplish downstream tasks, three components play a crucial role: a tokenizer, an input embedding layer, and module blocks. The tokenizer is responsible for converting the input text into the index representation that the model can understand, serving as a bridge between the input text and the model. The primary function of the input embedding layers is to transform the index representation into vector representations (also called embeddings). This process may involve techniques like word embeddings [52] and positional encodings [53] to capture vocabulary and positional information. These techniques enhance the input embedding layer’s ability to transform discrete data into meaningful continuous embeddings, facilitating subsequent processing. In the module blocks, the neural network structure is carefully designed to efficiently capture the complex relationships within embeddings. These blocks typically consist of multiple layers, with each layer responsible for processing different levels of features and abstract representations.
@@
Currently, LLMs can be mainly divided into three categories: autoregressive language models, masked language models, and encoder-decoder language models [26]. The autoregressive language models generate sequences based on the outputs from previous time steps. They are powerful for modeling the probability of sequences. In the code summarization task, the model recursively predicts each word to generate a code summary. The currently popular LLMs all belong to this category, such as GPT series [54], Codex [23], StarCoder [24], CodeGen [25], PolyCoder [26], and CodeLlama [55]. The masked language model can transform a given sequence into effective representations. They are trained using masked language modeling, a popular bidirectional objective function that aims to predict masked text pieces based on the surrounding context. Most of the early PLMs fall into this category, such as CodeBERT [56], UniXCoder [57], and GraphCodeBERT [58]. The encoder-decoder language models consist of a structure with two main components, i.e., encoder and decoder. The encoder is responsible for transforming the input sequence into an intermediate representation, while the decoder utilizes this intermediate representation to generate the output sequence. CodeT5 [29] and PLBART [59] are examples of such models in code. In recent times, the impact of LLMs has been truly transformative, revolutionizing both academic research and various industrial applications. In this paper, we focus on the application of autoregressive LLMs in software engineering, particularly in code summarization.
3 Methodology
3.1 Overview